OpenStudy (anonymous):

Find the slope of the line that contains the given pair of points: (-10,-15) and (2,-6)

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = (-6-(-15))/(2-(-10)) m = (-6+15)/(2+10) m = (9)/(12) m = 3/4 So the slope is 3/4

OpenStudy (anonymous):

thankyou

jimthompson5910 (jim_thompson5910):

yw